Grade 12th passMechanics

a pendulum is suspended vertically from a point. a force f in horizontal direction starts acting on it of constant magnitute ¾ of mg .
the minimum velocity required at the bottom to complete the vertical circle
/

Divyank Tyagi
8 Years agoGrade 12th pass

1 Answer

chhaya lohia
8 Years ago
Force acting =3/4mg which will be equal to the centripetal force required=mv^2/rSo equating both 3/4mg=mv^2/rSo v= under root 3gr/4